# A Comunication Framework for Compositional Generation

### v1

@@ Appendix A Recovering compositional structure from partially observed data
    what features of the pretrained encodings $\mathcal{G}\to\mathcal{D}$ and what requirements on the compositional structure of $\mathcal{G}$ would allow to actually guarantee this assumption,
@@
    alternatives for a rigorous definition of what is meant by ”to good accuracy” in the statement of assumption (A).
Once the above points are settled and defined, we can replace (A) by an explicit requirement. Since however these points seem like an ambitious whole research line and are far beyond the scope of this work, we leave the definition of ”to good accuracy” and the proof of reconstruction to future works.
Thus, introducing (A) allows us to prove fully rigorous statements such as Thm. A.3, in absence of a full theory of compositional reconstructon.
@@
  • It seems natural to conjecture that if assumption (A) does not hold for large value of $k$, then there is no hope for the compositional reconstruction to be achievable. Thus lowering the requirement on ”good enough accuracy” to make (A) hold seems the only avenue of future research.
Theorem A.1 is a consequence of below Theorem A.3 that proves that assumption (A) holds with probability tending to $1$ as $|\mathcal{G}|\to\infty$ and $|\mathcal{D}_{train}|/|\mathcal{G}|=p$ is bounded away from zero. What Theorem A.3 says is that for large sets of generators, it will become overwhelmingly unlikely that $\mathcal{D}_{train}$ does not contain enough data to reconstruct the generators.
A.1 Observing $\mathcal{D}_{train}$ allows to observe finite statistics of the factors with high probability
Our first result for this section is that when $\mathcal{D}_{train}\subseteq\mathcal{D}$ is chosen uniformly at random amongst sets of size $p|\mathcal{D}|$, for $|\mathcal{D}|=|\mathcal{G}|$ large enough, we find that with overwhelming probability all possible values of each of the factors $G_{1},\dots,G_{n}$ are achieved by some elements of $\mathcal{D}_{train}$. We state the result in Proposition A.2, whose proof in our view contains the main ideas useful also for the full result of Theorem A.3, which can be considered our main technical result for this section.
@@
$\mathbb{P}\{(\exists i\leq n)(\exists j\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>G_{i}\neq j\}$$\leq$ $nN\frac{(pN^{n})!((1-p)N^{n})!}{(N^{n})!}\frac{(N^{n}\tfrac{N-1}{N})!}{(pN^{n})!(N^{n}(\tfrac{N-1}{N}-p))!}$
@@
In fact Prop. A.2 is gives us guarantees that it will be possible to observe all options for each $G_{i}$ in the training set, however this may not be sufficient for the main aim of recovering the compositional structure of $\mathcal{G}$. However we next prove, in an extension of Prop. A.2, that actually we will also observe with high probability any possible combination of finitely many factors:
Theorem A.3.
Under the same hypotheses on $\mathcal{G},\mathsf{GenX},\mathcal{G}_{train},p$ as in Prop. A.2, for any $k\leq n$ we have the following.
$\mathbb{P}\{(\exists i_{1}<\dots<i_{k}\leq n)(\exists j_{1},\dots,j_{k}\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\}$
@@
As mentioned before, roughly the overall proof strategy is the same as for Prop. A.2. We introduce the sets $A^{j_{1},\dots,j_{k}}_{i_{1},\dots,i_{k}}=A^{\mathbf{j}}_{\mathbf{i}}$ and observe that the number of choices of $i_{1}<\cdots<i_{k}\leq n$ is $\tfrac{n!}{k!(n-k)!}$ while the number of choices of $j_{1},\dots,j_{k}\leq N$ is $N^{k}$. Thus the number of sets of the form $A^{\mathbf{j}}_{\mathbf{i}}$ is $N_{A}:=\tfrac{N^{k}n!}{k!(n-k)!}$. Furthermore, each set $A^{\mathbf{j}}_{\mathbf{i}}$ only imposes the constraint of entries of indices $i_{1},\dots,i_{k}$ from a $\mathbb{G}$ to avoid one combination of values $\vec{j}$, and thus it has cardinality $|A^{\mathbf{j}}_{\mathbf{i}}|=N^{n-k}(N^{k}-1)$. Then the number of choices of $\mathcal{G}_{train}\subseteq A^{\mathbf{j}}_{\mathbf{i}}$ of cardinality $pN^{n}$ is just the combinatorial coefficient $\tfrac{(N^{n-k}(N^{k}-1))!}{(pN^{n})!(N^{n-k}(N^{k}-1-pN^{k}))!}$ and by union bound we then find by the same reasoning as in Prop. A.2,
@@
This proves (A.3). In order to show that the bound tends to zero as $N^{n}\to\infty$ at fixed $p>0,k\leq n$, we first note that $\tfrac{n!}{k!(n-k)!}\leq n^{k}/k!$ and thus (as $k$ is assumed fixed) it suffices to show that
@@
Here again, we can proceed with the discussion of the two cases ($N$ bounded and $n\to\infty$, or $N\to\infty$) exactly as in the end of the proof of Prop. A.2, to conclude. ∎
Remark A.4 (Extension of Prop. A.2 and Theorems A.3 and A.1 to general $\mathcal{G}$).
If, for a choice of $n\geq 2$ and $N_{1},\dots,N_{n}\geq 2$ we have $\mathcal{G}=\{\mathbf{G}=[G_{1},\dots,G_{n}]:\ (\forall i\leq n)1\leq G_{i}\leq N_{i}\}$, then $|\mathcal{G}|=N_{1}\cdots N_{n}$ and by carefully following the same steps as in the proof of Prop. A.2 we find
$\mathbb{P}\{(\exists i\leq n)(\exists j\leq N_{i})(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>G_{i}\neq j\}\leq\sum_{i=1}^{n}N_{i}\left(1-\frac{1}{N}_{i}\right)^{pN_{1}\dots N_{n}},$ (7)
Again, this tends to zero for fixed $p$ as $|\mathcal{G}|=N_{1}\dots N_{n}\to\infty$: to prove it, we can proceed as follows. In a first case, if $n>2$ stays bounded, then we can discuss separately for each $1\leq i\leq n$ whether $N_{i}\to\infty$ or whether it stays bounded, as done in the proof of Prop. A.2, and in either case the corresponding summand on the left of (5) tends to zero, as in case (2) from the proof of Prop. A.2. Otherwise, $n\to\infty$ and we can proceed analogously to case (1) from the end of proof of Prop. A.2.
For the extension of Thm. A.3 to general $\mathcal{G}$ we get the bound
$\mathbb{P}\{(\exists i_{1}<\dots<i_{k}\leq n)(\exists j_{1},\dots,j_{k}\leq N)(\forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train})\>(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\}$
@@
Theorem A.1 is a direct consequence of assumption (A) and Theorem A.3, and thus it also directly extends to the general case.
Appendix B Regularization to counteract Holistic Encodings
In this section, we include rigorous results useful to justify the regularization of our Im phase, aimed at balancing NHE and EE.
This corresponds to the balance between compositional and parsimonious communication, a general theme in classical IL literature (see e.g. (Kirby & Hurford, 2002; Kirby et al., 2015)). It seems useful to specify more precise definitions in our generative setting, in order to help justify our architecture choices.
@@ Appendix B Regularization to counteract Holistic Encodings
    In §B.1 we set rigorous definitions of Holistic Encodings within the setting of §2.3 and §3.1.
@@
    In §B.2 we check that NHE is required for lowest Kolmogorov Complexity encodings.
@@
    We verify that optimizing for just UE and EE on our training set, will produce Holistic Encodings, and thus is not satisfactory for reconstructing $\mathcal{G}$ (see §B.3);
@@
    We give a few possible options on how to enrich the loss functions in our IL setup, aimed at establishing a better balance between EE and NHE (see §B.4).
B.1 Definitions of Holistic and Compositional encodings, and role of the NHE condition
In the classical IL setup (Kirby & Hurford, 2002; Kirby et al., 2015), a holistic system of communication is defined as an encoding $\mathcal{G}\to\mathcal{M}$ that maps $\mathcal{G}\ni\mathbf{G}\mapsto\mathbf{G^{\prime}}\in\mathcal{M}$ injectively but compl forgetting the compositional structure of the ”language” to be encoded, i.e. in our case, without respecting the structure of $\mathbf{G}$. In our generation task formalization, a difference is that the relevant encoding map now factors through the generation map as $\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}$. Now the generation map $\mathsf{GenX}$ is considered fixed, thus the definition refers to only $S$ and is the following:
Definition B.1.
In the setting from §2.3 and §3.1, sender $S:\mathcal{X}\to\mathcal{M}$ determines a holistic encoding if it satisfies UE but the messages associated to data generated as $\mathsf{GenX}(\mathbf{G})$ do not have a compositional structure in terms of generating factors $\mathbf{G}$, as specified by Definition B.2 below.
Note that, as in (Kirby & Hurford, 2002; Kirby et al., 2015), Definition B.1 of holistic encodings refers to ”respecting compositional structure”, which requires another definition. The definition of compositional encodings crucially depends on how complex we need/want to assume $\mathbf{G}$ to be. In this paper, for the sake of concreteness, we restricted to the simple case of $\mathbf{G}$ being formed as an element of a product space of finite cardinality, i.e., an $n$-ple of independent factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each $G_{i}$ can take finitely many values. We thus use a straightforward toy definition of compositional encodings, requiring that each of the $G_{i}$ is separately encoded as a string and these strings are concatenated together to give the message that encodes $\mathbf{G}$:
Definition B.2.

### v3

@@ Appendix D Recovering compositional structure from partially observed data
    what features of the pre-trained encodings $\mathcal{G}\to\mathcal{D}$ and what requirements on the compositional structure of $\mathcal{G}$ would allow to actually guarantee this assumption,
@@
    alternatives for a rigorous definition of what is meant by "to good accuracy" in the statement of assumption (A).
Once the above points are settled and defined, we can replace (A) by an explicit requirement. Since however these points seem like an ambitious whole research line and are far beyond the scope of this work, we leave the definition of "to good accuracy" and the proof of reconstruction to future works.
Thus, introducing (A) allows us to prove fully rigorous statements such as Thm. D.3, in absence of a full theory of compositional reconstruction.
@@
  • It seems natural to conjecture that if assumption (A) does not hold for large value of $k$, then there is no hope for the compositional reconstruction to be achievable. Thus lowering the requirement on "good enough accuracy" to make (A) hold seems the only avenue of future research.
Theorem D.1 is a consequence of below Theorem D.3 that proves that assumption (A) holds with probability tending to $1$ as $|\mathcal{G}|\to\infty$ and $|\mathcal{D}_{train}|/|\mathcal{G}|=p$ is bounded away from zero. What Theorem D.3 says is that for large sets of generators, it will become overwhelmingly unlikely that $\mathcal{D}_{train}$ does not contain enough data to reconstruct the generators.
D.1 Observing $\mathcal{D}_{train}$ allows to observe finite statistics of the factors with high probability
Our first result for this section is that when $\mathcal{D}_{train}\subseteq\mathcal{D}$ is chosen uniformly at random amongst sets of size $p|\mathcal{D}|$, for $|\mathcal{D}|=|\mathcal{G}|$ large enough, we find that with overwhelming probability all possible values of each of the factors $G_{1},\dots,G_{n}$ are achieved by some elements of $\mathcal{D}_{train}$. We state the result in Proposition D.2, whose proof in our view contains the main ideas useful also for the full result of Theorem D.3, which can be considered our main technical result for this section.
@@
$\mathbb{P}\left\{\begin{array}[]{l}\exists i\leq n,\ \exists j\leq N,\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train},\\ G_{i}\neq j\end{array}\right\}$$\leq$ $nN\frac{(pN^{n})!((1-p)N^{n})!}{(N^{n})!}\frac{(N^{n}\tfrac{N-1}{N})!}{(pN^{n})!(N^{n}(\tfrac{N-1}{N}-p))!}$
@@
In fact Prop. D.2 is gives us guarantees that it will be possible to observe all options for each $G_{i}$ in the training set, however this may not be sufficient for the main aim of recovering the compositional structure of $\mathcal{G}$. However we next prove, in an extension of Prop. D.2, that actually we will also observe with high probability any possible combination of finitely many factors:
Theorem D.3.
Under the same hypotheses on $\mathcal{G},\mathsf{GenX},\mathcal{G}_{train},p$ as in Prop. D.2, for any $k\leq n$ we have the following.
$\mathbb{P}\left\{\left.\begin{array}[]{l}\exists i_{1}<\dots<i_{k}\leq n\\ \exists j_{1},\dots,j_{k}\leq N\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train}\end{array}\right.(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\right\}$ (10)
@@
As mentioned before, roughly the overall proof strategy is the same as for Prop. D.2. We introduce the sets $A^{j_{1},\dots,j_{k}}_{i_{1},\dots,i_{k}}=A^{\mathbf{j}}_{\mathbf{i}}$ and observe that the number of choices of $i_{1}<\cdots<i_{k}\leq n$ is $\tfrac{n!}{k!(n-k)!}$ while the number of choices of $j_{1},\dots,j_{k}\leq N$ is $N^{k}$. Thus the number of sets of the form $A^{\mathbf{j}}_{\mathbf{i}}$ is $N_{A}:=\tfrac{N^{k}n!}{k!(n-k)!}$. Furthermore, each set $A^{\mathbf{j}}_{\mathbf{i}}$ only imposes the constraint of entries of indices $i_{1},\dots,i_{k}$ from a $\mathbb{G}$ to avoid one combination of values $\vec{j}$, and thus it has cardinality $|A^{\mathbf{j}}_{\mathbf{i}}|=N^{n-k}(N^{k}-1)$. Then the number of choices of $\mathcal{G}_{train}\subseteq A^{\mathbf{j}}_{\mathbf{i}}$ of cardinality $pN^{n}$ is just the combinatorial coefficient $\tfrac{(N^{n-k}(N^{k}-1))!}{(pN^{n})!(N^{n-k}(N^{k}-1-pN^{k}))!}$ and by union bound we then find by the same reasoning as in Prop. D.2,
@@
This proves (10). In order to show that the bound tends to zero as $N^{n}\to\infty$ at fixed $p>0,k\leq n$, we first note that $\tfrac{n!}{k!(n-k)!}\leq n^{k}/k!$ and thus (as $k$ is assumed fixed) it suffices to show that
@@
Here again, we can proceed with the discussion of the two cases ($N$ bounded and $n\to\infty$, or $N\to\infty$) exactly as in the end of the proof of Prop. D.2, to conclude. ∎
Remark D.4 (Extension of Prop. D.2 and Theorems D.3 and D.1 to general $\mathcal{G}$).
If, for a choice of $n\geq 2$ and $N_{1},\dots,N_{n}\geq 2$ we have $\mathcal{G}=\{\mathbf{G}=[G_{1},\dots,G_{n}]:\ (\forall i\leq n)1\leq G_{i}\leq N_{i}\}$, then $|\mathcal{G}|=N_{1}\cdots N_{n}$ and by carefully following the same steps as in the proof of Prop. D.2 we find
$\mathbb{P}\left\{\begin{array}[]{c}\exists i\leq n,\quad\exists j\leq N_{i},\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train},\\ \text{it holds }G_{i}\neq j\end{array}\right\}\leq\sum_{i=1}^{n}N_{i}\left(1-\frac{1}{N}_{i}\right)^{pN_{1}\dots N_{n}},$ (12)
Again, this tends to zero for fixed $p$ as $|\mathcal{G}|=N_{1}\dots N_{n}\to\infty$: to prove it, we can proceed as follows. In a first case, if $n>2$ stays bounded, then we can discuss separately for each $1\leq i\leq n$ whether $N_{i}\to\infty$ or whether it stays bounded, as done in the proof of Prop. D.2, and in either case the corresponding summand on the left of (5) tends to zero, as in case (2) from the proof of Prop. D.2. Otherwise, $n\to\infty$ and we can proceed analogously to case (1) from the end of proof of Prop. D.2.
For the extension of Thm. D.3 to general $\mathcal{G}$ we get the bound
$\mathbb{P}\left\{\left.\begin{array}[]{l}\exists i_{1}<\dots<i_{k}\leq n\\ \exists j_{1},\dots,j_{k}\leq N\\ \forall\mathbf{G}=[G_{1},\dots,G_{n}]\in\mathcal{G}_{train}\end{array}\right.(G_{i_{1}},\dots,G_{i_{k}})\neq(j_{1},\dots,j_{k})\right\}$ (16)
@@
Theorem D.1 is a direct consequence of assumption (A) and Theorem D.3, and thus it also directly extends to the general case.
Appendix E Proof of Prop. 3.1
We first recall the setting. Consider two fixed maps
$R^{\star}:V^{C}\to\mathcal{X},\quad S^{\star}:\mathcal{D}\subseteq\mathcal{X}\to V^{C},$
and let $\mathcal{R}_{[i]}:V^{C}\to V^{i}$ be the restriction to the first $i$ tokens, i.e. in previous notation, for $m=(m_{1},\dots,m_{C})\in V^{C}$ we set $\mathcal{R}_{[i]}(m):=m_{[i]}=(m_{1},\dots,m_{i})$.
For $R^{\star},S^{\star}$ as above and $S:\mathcal{D}\subseteq\mathcal{X}\to V^{C}$ we set, for $x\in\mathcal{D}$
$d_{i}^{S}(x):=d_{\mathcal{X}}(R^{\star}(S^{\star}(x)_{[i]})),R^{\star}(S(x)_{[i]}))=d_{\mathcal{X}}(R^{\star}(\mathcal{R}_{[i]}(S^{\star}(x)))),R^{\star}(\mathcal{R}_{[i]}(S(x)))).$
We also defined the following spaces:
$\mathcal{S}_{full}^{\epsilon}:=\Big{\{}S:\ \mathbb{E}_{x}\Big{[}\frac{1}{C}\sum_{i=1}^{C}d_{i}^{S}(x)\Big{]}<\epsilon\Big{\}},\quad\mathcal{S}_{final}^{\epsilon}:=\left\{S:\ \mathbb{E}_{x}\left[d_{C}^{S}(x)\right]<\epsilon\right\}.$ (18)
Then Prop. 3.1 states the following:
Proposition E.1.
With the above notations, we have the following
1. For $i<C$ and any choice of $M,\epsilon>0$, the restrictions on $i$-token submessages of $S(x)$ by requiring $S\in\mathcal{S}_{full}^{M}$ are more restrictive than those imposed by requiring $S\in\mathcal{S}_{final}^{\epsilon}$, i.e. we have $\mathcal{R}_{[i]}(\mathcal{S}_{full}^{M})\subset\mathcal{R}_{[i]}(\mathcal{S}_{final}^{\epsilon})$.
2. For the full message case $i=C$ in general we have $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{C\epsilon}$. If we further assume that $\mathbb{E}_{x}[d_{j}^{S}(x)]$ is non-increasing in $j$ then we have the stronger inclusion $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{\epsilon}$.
Proof.
Item 1 is direct, as no restriction on intermediate submessages $S(x)_{[i]}=\mathcal{R}_{[i]}(S(x))$ by the requirement $S\in\mathcal{S}_{final}^{\epsilon}$.
For item 2, we observe that
$\frac{1}{C}d_{C}^{S}(x)\leq\frac{1}{C}\sum_{i=1}^{C}d_{i}^{S}(x),$
and by taking expectation we find that $\mathbb{E}_{x}\left[\frac{1}{C}\sum_{i}d_{i}^{S}(x)\right]<\epsilon$ implies $\mathbb{E}_{x}[d_{C}^{S}(x)]\leq C\epsilon$, and thus $\mathcal{S}^{\epsilon}_{full}\subseteq\mathcal{S}^{C\epsilon}_{final}$, which gives the first part of the statement.
For the second part of item 2 we note that if $\mathbb{E}_{x}[d_{i}^{S}(x)]\geq\mathbb{E}_{x}[d_{C}^{S}(x)]$ for all $i\leq C$ we get
$\mathbb{E}_{x}\left[d_{C}^{S}(x)\right]\leq\min_{1\leq j\leq C}\mathbb{E}_{x}[d_{j}^{C}(x)]\leq\frac{1}{C}\sum_{i=1}^{C}\mathbb{E}_{x}\left[d_{i}^{S}(x)\right]=\mathbb{E}_{x}\left[\frac{1}{C}\sum_{i=1}^{C}d_{i}^{S}(x)\right],$
and thus $\mathcal{S}_{full}^{\epsilon}\subseteq\mathcal{S}_{final}^{\epsilon}$ as desired. ∎
Appendix F Regularization to counteract Holistic Encodings
In this section, we include rigorous results useful to justify the regularization of our Imitation phase, aimed at balancing compressibility and efficiency.
This corresponds to the balance between compositional and parsimonious communication, a general theme in classical IL literature (see e.g. [27, 29]). It seems useful to specify more precise definitions in our generative setting, in order to help justify our architecture choices.
@@ Appendix F Regularization to counteract Holistic Encodings
    In §F.1 we set rigorous definitions of Holistic Encodings within the setting of §2.1 and §2.2.
@@
    In §F.2 we check that compressibility is required for lowest Kolmogorov Complexity encodings.
@@
    We verify that optimizing for just expressivity and efficiency on our training set, will produce Holistic Encodings, and thus is not satisfactory for reconstructing $\mathcal{G}$ (see §F.3);
@@
    We give a few possible options on how to enrich the loss functions in our IL setup, aimed at establishing a better balance between efficiency and compressibility (see §F.4).
F.1 Definitions of Holistic and Compositional encodings, and role of the compressibility condition
In the classical IL setup [27, 29], a holistic system of communication is defined as an encoding $\mathcal{G}\to\mathcal{M}$ that maps $\mathcal{G}\ni\mathbf{G}\mapsto\mathbf{G^{\prime}}\in\mathcal{M}$ injectively but disregards the compositional structure of the "language" to be encoded, i.e. in our case, without respecting the structure of $\mathbf{G}$. In our generation task formalization, a difference is that the relevant encoding map now factors through the generation map as $\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}$. Now the generation map $\mathsf{GenX}$ is considered fixed, thus the definition refers to only $S$ and is the following:
Definition F.1.
In the setting from §2.1 and §2.2, sender $S:\mathcal{X}\to\mathcal{M}$ determines a holistic encoding if it satisfies expressivity but the messages associated to data generated as $\mathsf{GenX}(\mathbf{G})$ do not have a compositional structure in terms of generating factors $\mathbf{G}$, as specified by Definition F.2 below.
Note that, as in [27, 29], Definition F.1 of holistic encodings refers to "respecting compositional structure", which requires another definition. The definition of compositional encodings crucially depends on how complex we need/want to assume $\mathbf{G}$ to be. In this paper, for the sake of concreteness, we restricted to the simple case of $\mathbf{G}$ being formed as an element of a product space of finite cardinality, i.e., an $n$-ple of independent factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each $G_{i}$ can take finitely many values. We thus use a straightforward toy definition of compositional encodings, requiring that each of the $G_{i}$ is separately encoded as a string and these strings are concatenated together to give the message that encodes $\mathbf{G}$:
Definition F.2.